OpenStudy (anonymous):

Convert the standard form of the equation to slope intercept -4y=3-2x

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

slope intercept form y=mx+c

OpenStudy (gorv):

divid both side by -4

OpenStudy (gorv):

\[\frac{ -4*y }{ -4}=\frac{ 3 }{ -4 }-\frac{ 2x }{ -4 }\]

OpenStudy (gorv):

y=-3/4+2x/4 y=2x/4-3/4

OpenStudy (gorv):

y=x/2-3/4